Fermi National Laboratory
 

Operation Summary for December 9, 2009 - December 11, 2009

 

Wednesday Dec. 9

The day shift began with Operations monitoring the Tevatron (TeV) store 7422 (81.25E30) and the Recycler stash (221.6E10), with the Antiproton Source stacking (19.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

During the day shift, the Neutron Therapy Facility reported at 12:56 PM that they had completed treatment on their last patient for the day. TeV personnel began an end-of-store study at 1 PM. Operations terminated store 7422 at 3:07 PM. TeV RF work took place after the store was terminated.

During the evening shift, RF personnel tuned the TeV RF system. Operations put the TeV into shot setup and established store 7423 at 5:46 PM with an average initial luminosity (AIL) of 292.94E30. Unfortunately, the store was short-lived; it aborted (a clean abort with no quench) when the LCW temperature at TeV sector A2 climbed to 115˚ F and caused a TeV power supply to trip off. Operators called in Mechanical Support technicians to look at the sector A2 3-way valve. The techs rebuilt the A3 LCW pump (which was valved out late in the day shift due to seal leaks). Once the pump was rebuilt and brought back online, the 3-way valve regulation loops began to work properly. The Pelletron suffered 3 full discharges during this shift.

Thursday Dec. 10

The midnight shift began with Operations monitoring the stash (210.3E10), with Pbar stacking (28.9E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported several more full discharges of the Pelletron during shift. TeV Personnel began a Tevatron Electron Lens study early on the shift and concluded the study around 3:50 AM. Operations put the TeV into shot setup and established store 7427 at 5:10 AM with an AIL of 266.3E30. Operations reported downtime due to MI RF problems, which caused NuMI beam permit trips. Various temperature alarms popped up as the outside temp dropped into negative numbers.

During the day shift, Operations reported that store 7427 aborted at 9:50 AM due to a power failure at the TeV sector F1 cryo system building. The Duty Electrician reset the auxiliary service building breaker, but it tripped again at 11:49 AM. The Duty Electrician had the breaker changed out by 3:05 PM; recovery followed and lasted through to the end of the shift. Operations reported that the H- Source’s output was low all shift. Electron Cooling was down most of the shift.

During the evening shift, Operations reported that the TeV was recovering from the sector F1 cryo system power outage. The F1 cold compressor bypass valve air line was frozen again; Cryo techs thawed the line and had it operational by 5:45 PM. The F3 cold compressor failed again after F1 had recovered; Cryo techs had it repaired by 10:15 PM. A TeV quadrupole power supply (D0Q5) tripped with a ground fault indication - this was traced to LCW leaks on all the D0 low beta magnet quadrupole power supplies. The leaks were traced to the TeV sector C4 LCW system being off and the sector D1 3-way valve getting stuck open early in the shift. Mechanical Support techs had the C4 LCW system up and running by 10 PM, but EE Support techs required more time to finish cleaning up the mess at D0. Operations established a TeV ramp at 11:15 PM to warm up the TeV LCW system. Another TeV quadrupole power supply (C:B0Q6) tripped several times on a quench indication, but with little or no current. Operations called EE Support techs to check on the supply. The Pelletron suffered repeated full discharges during the shift – a stopgap running mode was implemented until an HEP store could be established. An access was planned to make repairs to the Pelletron and other machines after the next store gets established. A P2 transfer line vacuum burst traveled into the P1 and A1 lines and caused 1.5 hours of downtime. The Linac output remained low all shift.

Friday Dec. 11

The midnight shift began with Operations monitoring the stash (312.2E10), with Pbar stacking (10.4E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, the TeV LCW temperature climbed high enough that the D0 low beta supplies could be valved in and the hoses checked for additional leaks. Operations put the TeV into a dry squeeze while EE Support techs were investigating C:B0Q6 – they had this supply repaired by 4:30 AM. Operations put the TeV through dry and wet squeezes. Operations put the TeV into shot setup at 6:08 AM. Operations established store 7433 at 7:32 AM with an AIL of 240.3E30.

The Plans for Friday and the Weekend

The plans for today are to repair the Pelletron and conduct some necessary repairs in other machines. The plans for the weekend are to resume normal stacking, stashing, and storing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/14/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory